# Learning Smooth State-Dependent Traversability from Dense Point Clouds

### v1

@@ Abstract
A key open challenge in off-road autonomy is that the traversability of terrain often depends on the vehicle’s state. In particular, some obstacles are only traversable from some orientations. However, learning this interaction by encoding the angle of approach as a model input demands a large and diverse training dataset and is computationally inefficient during planning due to repeated model inference. To address these challenges, we present SPARTA, a method for estimating approach angle conditioned traversability from point clouds. Specifically, we impose geometric structure into our network by outputting a smooth analytical function over the 1-Sphere that predicts risk distribution for any angle of approach with minimal overhead and can be reused for subsequent queries. The function is composed of Fourier basis functions, which has important advantages for generalization due to their periodic nature and smoothness. We demonstrate SPARTA both in a high-fidelity simulation platform, where our model achieves a 91% success rate crossing a 40m boulder field (compared to 73% for the baseline), and on hardware, illustrating the generalization ability of the model to real-world settings.
@@ 1 Introduction
Figure 1: Dense point cloud (background) captures geometric detail of the environment (Top right), while discretizing dense point cloud into elevation map results in information loss. The underlying terrain is clearly distinguishable from noise (leaves, sensor noise, etc.) in the point cloud, but doing so is hard in an elevation map, leading to wrong risk estimation (Viridis Circle where Green denotes low risk).
Many applications for autonomous mobile robots involve traversing rough, off-road environments [1, 2, 3]. One key to reliably navigating in these environments is the ability to understand the interaction between the robot and the terrain, to assess relevant risk metrics (e.g., potential vehicle damage). Despite recent success in identifying traversable terrain [4, 5, 6, 7], motion planning in geometrically complex environments (e.g., dense forests, boulder fields) remains a challenge. For example, the risk associated with traversing a given terrain is often a complicated function of the robot’s state, especially its speed and angle of approach, and the terrain geometry111 We focus on angle-dependency in this paper, and leave the extension to other state variables to future work. . While high-speed traversal through risky terrain can be explicitly penalized within the planner, it is less clear how to account for the angle of approach (e.g., using hand-crafted risk heuristics based on terrain elevation [8, 9, 10]), as this requires both detailed terrain information and nuanced consideration of the underlying geometry.
@@
In this paper, we propose a technique to learn risk variables of interest, based on a point cloud representation of a given terrain patch. Specifically, to model system noise implicitly captured in the dataset, we represent risk as a categorical distribution due to its ease of use and capability of approximating multimodal distributions, as demonstrated in [4]. However, a point cloud-based network that naively encodes the angle of approach would be hard to train in practice due to the sparsity of data in robotics applications. As a result, our model takes in the point cloud and predicts an analytical function (represented using Fourier basis functions) that maps angle of approach to the target risk variable distribution. The choice of Fourier basis ensures that the resulting function is periodic, thus avoiding the wrap-around discontinuity [11], and has important advantages for generalization in the training process due to its upper-bounded Lipschitz constant by construction. The function, once computed for a terrain patch, can be queried efficiently with minimal computation overhead for any angle of approach at any timestep when integrated with modern planners like MPPI [12, 13]. To summarize, in this paper we propose SPARTA (Smooth Point-cloud Approach-angle Reasoning for Terrain Assessment) with the following key contributions:
@@
  • Evaluations of the proposed pipeline through learning to predict potential vehicle damage, a risk variable prior works handle using handcrafted rules, with extensive results both in a high-fidelity simulator [14] and on hardware, with integration into a MPPI planner [12].
@@ 2 Related Work
Traversability estimation algorithms aim to identify terrain suitable for robot navigation. Traversability estimates typically rely on either manually designed features [8, 10] or are derived from learned vehicle movement patterns [15, 4, 6, 7, 5]. Both [8] and [10] use local elevation measurements to estimate stepping difficulty using heuristics such as step-wise elevation change. On the other hand, learned methods usually rely on multiple sensing modalities, especially elevation maps and semantic maps [15, 4, 6, 7, 5]. WVN [5] uses camera input and trains a traversable terrain classifier online leveraging pretrained features from DINO-ViT [16]. Meanwhile, [6, 7] fuse semantic and elevation measurements, and use separate decoder heads to predict dense semantic and elevation maps, which are used to score planner rollouts using various heuristics. [15, 4] use semantic and elevation maps to predict traversability distributions represented as linear/angular traction and incorporate the worst-case traction into the vehicle dynamics model for risk-aware planning. However, none of these approaches models the angle-dependent nature of traversability. As a result, they can be overly conservative in geometrically complex environments, while underestimate traversal risk when the terrain is only dangerous from a few orientations.
@@
Unlike previous traversability estimation algorithms, SPARTA models the angle dependency of risk variables, which existing algorithms failed to capture, by exploiting the geometric structure of angle of approach (the 1-Sphere). Because we impose few assumptions on the risk variable of interest (i.e., it has clear lower and upper bounds and does not change abruptly with a small change in angle of approach), our approach could be used to extend (most) existing traversability estimation methods for more accurate and granular prediction. Besides tire deformation or potential vehicle damage demonstrated in our experiments (Section 4), we show a few other variables in off-road autonomy with a similar geometric structure in Appendix G.
@@ 6 Limitation
Our work has several limitations. For example, in this work we assume the environment is fully observable, i.e. a prebuilt point cloud map is accessible during planning. While the point cloud map we can build on the fly is sufficiently dense, the model’s performance without a prebuilt map is still to be tested. A potential solution to this problem is to penalize terrain with limited observation during planning or incorporate depth inpainting as in [7, 6]. In addition, this work focused on estimating traversability given geometric information, whereas semantic information is also important to consider in many off-road settings. We anticipate that future work could incorporate semantic information by performing sensor fusion with point pillar feature in birds-eye-view [7]. Last but not least, tire deformation is a metric that’s only available in simulation. However, we are not interested in predicting tire deformation in the real world, but rather using it as a surrogate metric for terrain property (i.e. how risky is the terrain). Future work could explore other variables to quantify vehicle damage.
@@ Appendix B Model Architecture
All the learned models described in this paper (Ours, AngleInput, and AngleFree) predict categorical distributions composed of $B=8$ bins, parameterized by the concentration parameters $[\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$.
@@ B.1 Neural Network Used in SPARTA
To represent the smooth analytical mapping $\phi\rightarrow\gamma_{\phi}^{i}$, we employ Fourier basis functions up to a maximum frequency of $n=3$, specifically the basis set $\{1,\cos(\phi),\sin(\phi),\cos(2\phi),\sin(2\phi),\cos(3\phi),\sin(3\phi)\}$. The choice of $n$ is justified in Appendix E with an ablation study. Consequently, for each bin $i\in[1,8]$, the network predicts 7 Fourier coefficients. Thus, the SPARTA decoder consists of a multilayer perceptron (MLP) that transforms the 256-dimensional point cloud features into a 56-dimensional output vector $[\mathbf{F}^{1},\dots,\mathbf{F}^{8}]$. This MLP decoder contains one hidden layer with 512 neurons and utilizes layer normalization and ReLU activation functions.
@@ B.2 AngleInput
The encoder for angle of approach $\phi$ is a MLP that maps 1-dimensional input $\phi$ to a 32-dimensional embedding with one hidden layer with 16 neurons. The decoder merges the point cloud feature (256-dimensional) and the angle feature (32-dimensional), and outputs the 8 concentration parameters $[\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with a hidden layer with 64 neurons. Both MLPs utilizes layer normalization and ReLU activation functions.
@@ B.3 AngleFree
The AngleFree baseline network uses the same convolutional backbone as AngleInput. A decoder head transforms the 256-dimensional point cloud feature to the 8 concentration parameters $[\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with two hidden layers with [128, 64] neurons.
@@ Appendix C Inference Efficiency
To demonstrate the query efficiency of our model compared to AngleInput, we measure their respective inference runtimes. Specifically, for AngleInput, we measure the runtime of encoding the approach angle, decoding the concentration parameters, and computing the categorical distribution (LABEL:eqn:PMF). For our model, we measure the runtime associated only with computing the concentration parameters (LABEL:eqn:fourier) and the categorical distribution (LABEL:eqn:PMF). Note other components (e.g. extracting feature vector from point cloud) of both models can be precomputed or reused. Considering that more complex point clouds require higher-dimensional point cloud features and deeper network architectures, we perform a stress test by progressively increasing the number of layers in both models, as well as the maximum frequency $n$ in our model. Both models process 25,000 queries with average runtimes reported in Fig. 3. Our model demonstrates an order-of-magnitude faster runtime compared to AngleInput. Most importantly, the runtime of our model remains constant despite increased complexity, as it only involves a single batched, low-dimensional dot product and sigmoid computation  [^2] The dimension of the dot product depends only on the choice of maximum frequency $n$. In fact the dimension is exactly $12n+$ 1.. In contrast, the runtime of AngleInput increases with complexity.
@@ Appendix F Hardware Specification
Appendix G Geometric Structure in Off-road Autonomy
@@
In this section, we demonstrate other problem settings that could benefit from modeling angle-dependency. We use the angle-dependency of linear traction (the robot’s ability to execute commanded velocity on the given terrain [4]) as an example, and show how our approach can be applied to this problem setting. To construct an empirical dataset, we command the vehicle to drive at a constant linear velocity of 0.5 m/s over the obstacle, along the forward (Blue) and the backward (Orange) direction shown in Fig. 6(a). We record the achieved linear velocity (estimated via DLIO [26]), and compute the linear traction as the ratio between the achieved linear velocity and the commanded linear velocity (0.5 m/s). The data points are discretized into histograms and then normalized into categorical distributions. As shown in Fig. 6(b), the backward traversal (Orange) achieved significantly lower linear tractions compared to the forward traversal (Blue). This matches our intuition because the obstacle is smooth from the forward direction and sharp from the other. A system that models the angle-dependency of linear traction would thus be able to capture this difference. Notice SPARTA can be applied to this problem setting with minimal changes (e.g. adjusting input point cloud size) using an empirical dataset of linear traction distributions. However, systems that do not model angle-dependency (Green) [4] could be overly conservative and never command the robot to traverse the obstacle from the forward direction, which is traversable in practice.

### v2

@@ Abstract
A key open challenge in off-road autonomy is that the traversability of terrain often depends on the vehicle’s state. In particular, some obstacles are only traversable from some orientations. However, learning this interaction by encoding the angle of approach as a model input demands a large and diverse training dataset and is computationally inefficient during planning due to repeated model inference. To address these challenges, we present SPARTA, a method for estimating approach angle conditioned traversability from point clouds. Specifically, we impose geometric structure into our network by outputting a smooth analytical function over the 1-Sphere that predicts risk distribution for any angle of approach with minimal overhead and can be reused for subsequent queries. The function is composed of Fourier basis functions, which has important advantages for generalization due to their periodic nature and smoothness. We demonstrate SPARTA both in a high-fidelity simulation platform, where our model achieves a 91% success rate crossing a 40m boulder field (compared to 73% for the baseline), and on hardware, illustrating the generalization ability of the model to real-world settings. Our code will be available at https://github.com/neu-autonomy/SPARTA.
@@ 1 Introduction
Figure 1: Dense point cloud captures geometric detail of the environment (Top right), while discretizing it into elevation map results in information loss. The underlying terrain is clearly distinguishable from noise (leaves, sensor noise) in the point cloud, but doing so is hard in an elevation map, leading to wrong risk estimation (Viridis Circle where Green denotes low risk).
Many applications for autonomous mobile robots involve traversing rough, off-road environments [1, 2, 3]. One key to reliably navigate in these environments is to estimate terrain traversability, i.e., the ability to understand the interaction between the robot and the terrain, and assess relevant risk metrics (e.g., potential vehicle damage). Despite recent success in identifying traversable terrain [4, 5, 6, 7], motion planning in geometrically complex environments (e.g., dense forests, boulder fields) remains a challenge. Specifically, the risk associated with traversing a given terrain is often a complicated function of the robot’s state, for example its speed and angle of approach, and the terrain geometry [^1] We focus on angle-dependency in this paper, and leave the extension to other state variables to future work. . It is especially unclear how to account for the angle of approach (e.g., using hand-crafted risk heuristics based on terrain elevation [8, 9, 10]), as this requires both detailed terrain information and nuanced consideration of the underlying geometry.
@@
In this paper, we propose a technique to learn risk variables of interest, based on a point cloud representation of the terrain. Specifically, to model system noise implicitly captured in the dataset, we represent risk as a categorical distribution due to its ability to approximate arbitrary distributions [4]. However, a point cloud-based network that naively encodes the angle of approach would be hard to train in practice due to the sparsity of data in robotics applications. As a result, our model takes in the point cloud and predicts an analytical function (represented using Fourier basis functions) that maps angle of approach to the target risk variable distribution. The choice of Fourier basis ensures that the resulting function is periodic, thus avoiding the wrap-around discontinuity [11], and has important advantages for generalization in the training process due to its upper-bounded Lipschitz constant by construction. The function, once computed for a terrain patch, can be queried efficiently with minimal computation overhead for any angle of approach at any timestep when integrated with modern planners like MPPI [12, 13]. To summarize, in this paper we propose SPARTA (Smooth Point-cloud Approach-angle Reasoning for Terrain Assessment) with the following key contributions:
@@
  • Evaluations of the proposed pipeline through learning to predict potential vehicle damage, a risk variable prior works handle using handcrafted rules, with extensive results both in a high-fidelity simulator [14] and on hardware, with integration into a MPPI planner [12]. In Appendix G we discuss SPARTA’s generalization to other problem settings.
@@ 2 Related Work
Traversability estimation aims to identify terrain suitable for robot navigation. Traversability estimates typically rely on either manually designed features [8, 10] or are derived from learned vehicle movement patterns [15, 4, 6, 7, 5]. Both [8] and [10] use local elevation measurements to estimate stepping difficulty using heuristics such as step-wise elevation change. On the other hand, learned methods usually rely on multiple sensing modalities, especially elevation maps and semantic maps [15, 4, 6, 7, 5]. WVN [5] uses camera input and trains a traversable terrain classifier online leveraging pretrained features from DINO-ViT [16]. Meanwhile, [6, 7] fuse semantic and elevation measurements, and use separate decoder heads to predict dense semantic and elevation maps, which are used to score planner rollouts using various heuristics. [15, 4] use semantic and elevation maps to predict traversability distributions represented as linear/angular traction and incorporate the worst-case traction into the vehicle dynamics model for risk-aware planning. Several works modeled the state-dependency of traversability by encoding robot state as part of the neural networks’ input [17, 18, 19]. However, there performances suffer from data sparsity and the dimensionality mismatch between exteroceptive inputs and low dimensional state space.
@@
Unlike previous traversability estimation algorithms, SPARTA explicitly models the angle dependency of risk variables by exploiting the geometric structure of angle of approach (the 1-Sphere). Because we impose few assumptions on the risk variable of interest (i.e., it has clear lower and upper bounds and does not change abruptly with a small change in angle of approach), our approach could be used to extend (most) existing traversability estimation methods for more accurate and granular prediction. Besides tire deformation (demonstrated in our experiments Section 4), we show examples of variables with a similar geometric structure in Appendix G.
@@ 6 Limitation
Our work has several limitations. For example, in this work we assume the environment is fully observable, i.e. a prebuilt point cloud map is accessible during planning. While the point cloud map we can build on the fly is sufficiently dense, the model’s performance without a prebuilt map is still to be tested. A potential solution to this problem is to penalize terrain with limited observation during planning or incorporate depth inpainting as in [7, 6]. In addition, this work focused on estimating traversability given geometric information, whereas semantic information is also important to consider in many off-road settings. We anticipate that future work could incorporate semantic information by performing sensor fusion with point pillar feature in birds-eye-view [7]. While the smoothness assumption in Section 3.3 is reasonable, settings that potentially break this assumption might exist in the real world. Quantitative analysis to identify such failure modes can be a valuable addition to our framework. Last but not least, tire deformation is a metric that’s only available in simulation. However, we are not interested in predicting tire deformation in the real world, but rather using it as a surrogate metric for terrain property (i.e. how risky is the terrain). Future work could explore other variables to quantify vehicle damage.
@@ Appendix B Model Architecture
All the learned models described in this paper (Ours, AngleInput, and AngleFree) predict categorical distributions composed of $B=8$ bins, parameterized by the concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$.
@@ B.1 Neural Network Used in SPARTA
To represent the smooth analytical mapping $\phi\rightarrow\gamma_{\phi}^{i}$, we employ Fourier basis functions up to a maximum frequency of $n=3$, specifically the basis set $\{1,\cos(\phi),\sin(\phi),\cos(2\phi),\sin(2\phi),\cos(3\phi),\sin(3\phi)\}$. The choice of $n$ is justified in Appendix E with an ablation study. Consequently, for each bin $i\in[1,8]$, the network predicts 7 Fourier coefficients. Thus, the SPARTA decoder consists of a multilayer perceptron (MLP) that transforms the 256-dimensional point cloud features into a 56-dimensional output vector $\mathbf{F}^{1},\dots,\mathbf{F}^{8}]$. This MLP decoder contains one hidden layer with 512 neurons and utilizes layer normalization and ReLU activation functions.
@@ B.2 AngleInput
The encoder for angle of approach $\phi$ is a MLP that maps 1-dimensional input $\phi$ to a 32-dimensional embedding with one hidden layer with 16 neurons. The decoder merges the point cloud feature (256-dimensional) and the angle feature (32-dimensional), and outputs the 8 concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with a hidden layer with 64 neurons. Both MLPs utilizes layer normalization and ReLU activation functions.
@@ B.3 AngleFree
The AngleFree baseline network uses the same convolutional backbone as AngleInput. A decoder head transforms the 256-dimensional point cloud feature to the 8 concentration parameters $\gamma_{\phi}^{1},\dots,\gamma_{\phi}^{8}]$, with two hidden layers with [128, 64] neurons.
@@ Appendix C Inference Efficiency
To demonstrate the query efficiency of our model compared to AngleInput, we measure their respective inference runtimes. Specifically, for AngleInput, we measure the runtime of encoding the approach angle, decoding the concentration parameters, and computing the categorical distribution (LABEL:eqn:PMF). For our model, we measure the runtime associated only with computing the concentration parameters (LABEL:eqn:fourier) and the categorical distribution (LABEL:eqn:PMF). Note other components (e.g. extracting feature vector from point cloud) of both models can be precomputed or reused. Considering that more complex point clouds require higher-dimensional point cloud features and deeper network architectures, we perform a stress test by progressively increasing the number of layers in both models, as well as the maximum frequency $n$ in our model. Both models process 25,000 queries with average runtimes reported in Fig. 3. Our model demonstrates an order-of-magnitude faster runtime compared to AngleInput. Most importantly, the runtime of our model remains constant despite increased complexity, as it only involves a single batched, low-dimensional dot product and sigmoid computation  [^2] The dimension of the dot product depends only on the choice of maximum frequency $n$. In fact the dimension is exactly $2n+1$.. In contrast, the runtime of AngleInput increases with complexity.
@@ Appendix F Hardware Specification
Appendix G Generalization to Other Problem Settings
@@
In this section, we first demonstrate other problem settings that could benefit from modeling angle-dependency. We use the angle-dependency of linear traction (the robot’s ability to execute commanded velocity on the given terrain [4]) as an example, and show how our approach can be applied to this problem setting. To construct an empirical dataset, we command the vehicle to drive at a constant linear velocity of 0.5 m/s over the obstacle, along the forward (Blue) and the backward (Orange) direction shown in Fig. 6(a). We record the achieved linear velocity (estimated via DLIO [29]), and compute the linear traction as the ratio between the achieved linear velocity and the commanded linear velocity (0.5 m/s). The data points are discretized into histograms and then normalized into categorical distributions. As shown in Fig. 6(b), the backward traversal (Orange) achieved significantly lower linear tractions compared to the forward traversal (Blue). This matches our intuition because the obstacle is smooth from the forward direction and sharp from the other. A system that models the angle-dependency of linear traction would thus be able to capture this difference. Notice SPARTA can be applied to this problem setting with minimal changes (e.g. adjusting input point cloud size) using an empirical dataset of linear traction distributions. However, systems that do not model angle-dependency (Green) [4] could be overly conservative and never command the robot to traverse the obstacle from the forward direction, which is traversable in practice.
In fact, we anticipate any variables that exhibit smooth and continuous behaviors, especially those that depends on variables with Lie group structure, could similarly benefit from our representation framework. SPARTA can be easily modified to use their respective basis functions and predict the corresponding coefficients.